A property was purchased for $9330.00 down and payments of $1373.00 at the end of every month for 3 years. Interest is 11% per annum compounded annually. What was the purchase price of the property? How much is the cost of financing?

The purchase price of the property was $____?

The cost of financing is $______?


pleaseeee help !!

!!!!

¡¡¡
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
11 = ((1+Stated rate%/(12*100))^12-1)*100
Stated rate% = 10.4815
PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 1373*((1-(1+ 10.4815/1200)^(-3*12))/(10.4815/1200))
PV = 42254.32

Property value = PV+down = 42254.32+9330 =

51584.32

Cost of financing = monthly payments*number of months-PV = 1373*12*3-42254.32

=

7173.68